# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v1

@@ 4.1. Seiberg-Witten invariants of smoothly reducible double coverings
The following proposition is a major trait in our proof of Theorem D. We direct the reader towards [31, Section 2.4], [8, Section 2.5] and the references there for the necessary background on Seiberg-Witten theory.
@@
(Kotschick-Morgan-Taubes [25, Proposition 2], Baykur-Stipsicz-Szabó [8, Corollary 2.11]). Let $X_{1}$ and $X_{2}$ be closed oriented smooth four-manifolds and suppose that $b_{1}(X_{2})=0=b_{2}^{+}(X_{2})$. There is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s}_{X})$ on the connected sum $X=X_{1}\mathbin{\#}X_{2}$ that extends a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$ such that
@@
Baykur-Stipsicz-Szabó observed in [8, Corollary 2.11] that Proposition 17 is a special case of the Seiberg-Witten blow-up formula [12]. Proposition 18, along with the blow up formula [12] (see [8, Proposition 2.8]) will be used to argue the conclusion of the irreduciblity property in the statement of Theorem D, together with an argument due to Baykur-Stipsicz-Szabó [8].
@@
Let $\widetilde{M}$ be a closed smooth oriented simply connected four-manifold with exactly two Seiberg-Witten basic classes ±$\pm K_{\widetilde{M}}\in\mathfrak{B}_{\widetilde{X}}$. Suppose there is an orientation-preserving fixed-point-free involution
@@
Consider the map
@@
assigning to a given $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s})$ its first Chern class with $\mathbb{Q}$-coefficients. As argued by Baykur-Stipsicz-Szabó [8, Proof of Proposition 6.2], the blow up formula [12] and Proposition 17 imply that the image of the set
@@
under the map (4.4) has two elements, given that the four-manifold $\widetilde{M}$ is assumed to have exactly two Seiberg-Witten basic classes and the covering decomposes as $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$ by Lemma 7. On the other hand, the decomposition (4.3) implies that the image of the set (4.5) under (4.4) contains at least four elements. This is a contradiction and we conclude that the four-manifold $M(2p)$ is irreducible. The claim on the value (4.1) of the Seiberg-Witten invariant is an immediate consequence of Proposition 17.
@@ 4.2. Proof of Theorem D
Denote by $X_{n}(2)\mathrel{\mathop{:}}=\widetilde{X}_{n}/\tau_{n}$ the closed smooth oriented four-manifold with fundamental group of order two that arises as the quotient of the closed oriented simply connected four-manifold $\widetilde{X}_{n}$ by its orientation-preserving fixed-point free involution $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$. The classification result of Hambleton-Kreck [20, Theorem C] says that $X_{n}(2)$ is homeomorphic to $Y\mathbin{\#}\Sigma_{2,k}$ for a given closed simply connected topological four-manifold $Y$ with Kirby-Siebenmann invariant $\operatorname{KS}(Y)=k=\Sigma_{2,k}$. Fix an odd positive integer $p\in\mathbb{Z}_{>0}$ and assemble
@@
and $S_{p}$ is the compact four-manifold that was manufactured in Section 2.1. Lemma 6 says that the fundamental group of the four-manifold (4.6) is the finite cyclic group $\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$. Moreover, there is a homeomorphism between $X_{n}(2p)$ and $Y\mathbin{\#}\Sigma_{2p,k}$ by Corollary 5 (and [20, Theorem C] as well).
Lemma 7 says that the degree-two covering $X_{n}(p)$ that corresponds to the subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to $X_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}$. Proposition 18 says that the four-manifold (4.6) is irreducible. To conclude on the existence of infinitely many pairwise non-diffeomorphic irreducible smooth structures on $Y\mathbin{\#}\Sigma_{2p,k}$, we employ Lemma 8 and distinguish them in terms of their double coverings: the value of the Seiberg-Witten function is assumed to be $\operatorname{SW}_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$ on its sole pair of basic classes and $\widetilde{X}_{n_{1}}$ is diffeomorphic to $\widetilde{X}_{n_{2}}$ if and only if $n_{1}=n_{2}$ (cf. [31], [8, Proposition 2.14]). $\square$
@@
A substantial region of Theorem 19 is due to Arabadji-Porter’s extensive study on the existence of fixed-point free orientation-preserving involutions on the closed irreducible symplectic simply connected four-manifolds constructed in [2, 18]. A thorough study of the irreducibility of these simply connected four-manifolds guarantees that they satisfy the hypothesis of Theorem D. $\square$
@@ 6.4. $\operatorname{Pin}^{+}$-cobordism classes and a computational tool
$\bullet$ Kirby-Taylor [24, Theorem 5.2]. There is a group isomorphism
@@
[Proof]We first assemble the $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$ by gluing together $\operatorname{Pin}^{+}$-structures $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ and $(N_{p},\phi_{N_{p}})$ that induce the same $\operatorname{Pin}^{+}$-structure on their common boundary; see (6.3). The $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ induces a $\operatorname{Pin}^{+}$-structure $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ as a codimension zero submanifold [24, Construction 1.5], which restricts to a $\operatorname{Pin}^{+}$-structure
@@
there are two choices of such structures [32, Corollary 6.4] and there is a one-to-one correspondence between the initial $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ and the $\operatorname{Pin}^{+}$-structure (6.6) (see [23, Proposition 1 Chapter IV]). We now argue that there is a $\operatorname{Pin}^{+}$-structure $(N_{p},\phi_{N_{p}})$ that induces the $\operatorname{Pin}^{+}$-structure (6.6) on $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Since $w_{2}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=0$ and $H^{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z}/2)=\mathbb{Z}/2$, there are two $\operatorname{Pin}^{+}$-structures on $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ [32, Corollary 6.4]. The hypothesis $p\neq 0\mod 2$ allows us to pick the $\operatorname{Pin}^{+}$-structure $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ that induces $(N_{p},\phi_{N_{p}})$ such that the induced $\operatorname{Pin}^{+}$-structure $(\partial N_{p},\phi_{N_{p}}|_{\partial})$ matches (6.6). We conclude that there is a $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$. Next, we put together the compact five-manifold $V$ by gluing the product cobordisms
@@ 6.5. Proof of Theorem F
and it follows that there are two $\operatorname{Pin}^{+}$-structures $(A(2p),\pm\phi_{A(2p)})$ and two $\operatorname{Pin}^{+}$-structures $(B(2p),\pm\phi_{B(2p)})$ [32, Corollary 6.4], . Moreover, the $\operatorname{Pin}^{+}$-structures on the same underlying smooth four-manifold are mutual inverses in the fourth $\operatorname{Pin}^{+}$-cobordism group $\Omega^{\operatorname{Pin}^{+}}_{4}$ and the values
@@
belong to different classes in $\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16$ for any $k\in\mathbb{Z}_{>0}$ and that the four-manifolds $A(2p)$ and $B(2p)$ are not stably-diffeomorphic. This settles the claim of Item (5).
Item (6) follows the third clause of Theorem 25 and the corresponding coveringings of the decompositions (6.7) and (6.8). $\square$

### v2

@@ 4.1. Seiberg-Witten invariants of smoothly reducible double coverings
The following proposition is a major trait in our proof of Theorem D. We direct the reader towards [34, Section 2.4], [8, Section 2.5] and the references there for the necessary background on Seiberg-Witten theory.
@@
(Kotschick-Morgan-Taubes [27, Proposition 2], Baykur-Stipsicz-Szabó [8, Corollary 2.11]). Let $X_{1}$ and $X_{2}$ be closed oriented smooth four-manifolds and suppose that $b_{1}(X_{2})=0=b_{2}^{+}(X_{2})$. There is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s}_{X})$ on the connected sum $X=X_{1}\mathbin{\#}X_{2}$ that extends a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$ such that
@@
Baykur-Stipsicz-Szabó observed in [8, Corollary 2.11] that Proposition 18 is a special case of the Seiberg-Witten blow-up formula [13]. Proposition 19, along with the blow up formula [13] (see [8, Proposition 2.8]). Proposition 18 will be used to argue the conclusion of the irreduciblity property in the statement of Theorem D, together with an argument due to Baykur-Stipsicz-Szabó [8].
@@
(cf. [8, Proof of Proposition 6.2]). Let $\widetilde{M}$ be a closed smooth oriented simply connected four-manifold and assume it satisfies at least one of the following two properties.
$\bullet$ The four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes ±$\pm K_{\widetilde{M}}\in\mathfrak{B}_{\widetilde{M}}$ or
$\bullet$ The four-manifold $\widetilde{M}$ is minimal and it has several Seiberg-Witten basic classes {±$\{\pm K^{1}_{\widetilde{M}},\ldots,\pm K^{i}_{\widetilde{M}}\}\in\mathfrak{B}_{\widetilde{M}}$ with $i\geq 2$.
Suppose there is an orientation-preserving fixed-point-free involution
@@
The reader might have already noticed that the properties that the simply connected four-manifold $\widetilde{M}$ of Proposition 19 is required to have imply that it is irreducible.
@@
At this point, we employ an argument due to Baykur-Stipsicz-Szabo [8] to prove Proposition 19. Consider the map
@@
assigning to a given $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s})$ its first Chern class with $\mathbb{Q}$-coefficients. Suppose first that the four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes. As argued by Baykur-Stipsicz-Szabó [8, Proof of Proposition 6.2], the blow up formula [13] (see [8, Proposition 2.8]) and Proposition 18 imply that the image of the set
@@
under the map (4.4) has two elements, given that the four-manifold $\widetilde{M}$ is assumed to have exactly two Seiberg-Witten basic classes and the covering decomposes as $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$ by Lemma 8. On the other hand, the decomposition (4.3) implies that the image of the set (4.5) under (4.4) contains at least four elements. This is a contradiction and we conclude that the four-manifold $M(2p)$ is irreducible. A slight extension of Baykur-Stipsicz-Szabó’s argument allows us to reach a similar contradiction and conclude that $M(2p)$ is irreducible under the assumptions that the four-manifold $\widetilde{M}$ admits several Seiberg-Witten basic classes and that it is minimal. In this case, the image of the set (4.5) under the map (4.4) has $i$ elements, where $i>2$ is the number of Seiberg-Witten basic classes of $\widetilde{M}$. If there were indeed a decomposition as in (4.3), then the number of elements of the image of the set (4.5) under the map (4.4) would be strictly larger than $i$. We conclude that the four-manifold $M(2p)$ is irreducible. The claim on the value (4.1) of the Seiberg-Witten invariant is an immediate consequence of Proposition 18.
@@ 4.2. Proof of Theorem D
Denote by $X_{n}(2)\mathrel{\mathop{:}}=\widetilde{X}_{n}/\tau_{n}$ the closed smooth oriented four-manifold with fundamental group of order two that arises as the quotient of the closed oriented simply connected four-manifold $\widetilde{X}_{n}$ by its orientation-preserving fixed-point free involution $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$. The classification result of Hambleton-Kreck [21, Theorem C] says that $X_{n}(2)$ is homeomorphic to $Y\mathbin{\#}\Sigma_{2,k}$ for a given closed simply connected topological four-manifold $Y$ with Kirby-Siebenmann invariant $\operatorname{KS}(Y)=k=\Sigma_{2,k}$. Fix an odd positive integer $p\in\mathbb{Z}_{>0}$ and assemble
@@
and $S_{p}$ is the compact four-manifold that was manufactured in Section 2.1. Lemma 6 says that the fundamental group of the four-manifold (4.6) is the finite cyclic group $\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$. Moreover, there is a homeomorphism between $X_{n}(2p)$ and $Y\mathbin{\#}\Sigma_{2p,k}$ by Corollary 5 (and [21, Theorem C] as well).
Lemma 8 says that the degree-two covering $X_{n}(p)$ that corresponds to the subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to $X_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}$. Proposition 19 says that the four-manifold (4.6) is irreducible. At this point, we have proven all clauses of Theorem D except for the third one. To conclude on the existence of infinitely many pairwise non-diffeomorphic irreducible smooth structures on $Y\mathbin{\#}\Sigma_{2p,k}$ and conclude the proof of the theorem, we employ Lemma 9 and distinguish them in terms of their double coverings: the value of the Seiberg-Witten function is assumed to be $\operatorname{SW}_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$ on its sole pair of basic classes and $\widetilde{X}_{n_{1}}$ is diffeomorphic to $\widetilde{X}_{n_{2}}$ if and only if $n_{1}=n_{2}$. Notice that this argument is due to Stipsicz-Szabó [34] and Baykur-Stipsicz-Szabó [8, Proposition 2.14]. $\square$
@@
A substantial region of Theorem 20 is due to Arabadji-Morgan’s extensive study on the existence of fixed-point free orientation-preserving involutions on the closed irreducible symplectic simply connected four-manifolds constructed in [2, 19]. Theorem D follows from Theorem 20 and the last clause of Theorem D. $\square$
Remark 21.
As Porter Morgan kindly pointed out to us, several of the irreducible simply-connected four-manifolds with an involution constructed in [4] have more than two Seiberg-Witten basic classes. A stronger conclusion of infinitely many exotic smooth structures in Theorem E using our mechanism would depend on an improvement of Arabadji-Morgan’s result in this direction.
Remark 22.
It is interesting to understand what is obtained from applying our mechanism to other constructions of exotic four-manifolds like the recent ones due to Davis-Hayden-Huang-Ruberman-Sunukjian [11] and to Harris-Naylor-Park [23], for example.
@@ 6.4. $\operatorname{Pin}^{+}$-cobordism classes and a computational tool
$\bullet$ Kirby-Taylor [26, Theorem 5.2]. There is a group isomorphism
@@
[Proof]We first assemble the $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$ by gluing together $\operatorname{Pin}^{+}$-structures $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ and $(N_{p},\phi_{N_{p}})$ that induce the same $\operatorname{Pin}^{+}$-structure on their common boundary; see (6.3). The $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ induces a $\operatorname{Pin}^{+}$-structure $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ as a codimension zero submanifold [26, Construction 1.5], which restricts to a $\operatorname{Pin}^{+}$-structure
@@
there are two choices of such structures [35, Corollary 6.4] and there is a one-to-one correspondence between the initial $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ and the $\operatorname{Pin}^{+}$-structure (6.6) (see [25, Proposition 1 Chapter IV]). We now argue that there is a $\operatorname{Pin}^{+}$-structure $(N_{p},\phi_{N_{p}})$ that induces the $\operatorname{Pin}^{+}$-structure (6.6) on $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Since $w_{2}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=0$ and $H^{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z}/2)=\mathbb{Z}/2$, there are two $\operatorname{Pin}^{+}$-structures on $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ [35, Corollary 6.4]. The hypothesis $p\neq 0\mod 2$ allows us to pick the $\operatorname{Pin}^{+}$-structure $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ that induces $(N_{p},\phi_{N_{p}})$ such that the induced $\operatorname{Pin}^{+}$-structure $(\partial N_{p},\phi_{N_{p}}|_{\partial})$ matches (6.6). We conclude that there is a $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$. Next, we put together the compact five-manifold $V$ by gluing the product cobordisms
@@ 6.5. Proof of Theorem F
and it follows that there are two $\operatorname{Pin}^{+}$-structures $(A(2p),\pm\phi_{A(2p)})$ and two $\operatorname{Pin}^{+}$-structures $(B(2p),\pm\phi_{B(2p)})$ [35, Corollary 6.4], . Moreover, the $\operatorname{Pin}^{+}$-structures on the same underlying smooth four-manifold are mutual inverses in the fourth $\operatorname{Pin}^{+}$-cobordism group $\Omega^{\operatorname{Pin}^{+}}_{4}$ and the values
@@
belong to different classes in $\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16$ for any $k\in\mathbb{Z}_{>0}$ and that the four-manifolds $A(2p)$ and $B(2p)$ are not stably-diffeomorphic. This settles the claim of Item (5). Item (6) follows from the third clause of Theorem 28 and the corresponding coverings of the decompositions (6.7) and (6.8).
Finally, to construct the closed non-smoothable topological four-manifold $\ast A(2p)$, we substitute $\mathbb{R}P^{4}$ with the closed non-smootable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space in (6.7). The latter four-manifold exists by a computation using the surgery exact sequence [39, Chapter 14] and an explicit construction of it was given by Ruberman in [32]. The existence and uniqueness results of a topological tubular neighborhood $\nu(\alpha_{\ast\mathbb{R}P^{4}})$ for a locally flat simple loop $\alpha_{\ast\mathbb{R}P^{4}}\subset\ast\mathbb{R}P^{4}$ are found in [16, Section 5.1]. Choose a loop that generates the fundamental group and assemble
(6.11) $\ast A(2p)\mathrel{\mathop{:}}=(\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\setminus\nu(\alpha_{\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}}))\cup N_{p}.$
The existence of a homotopy equivalence $\ast\mathbb{R}P^{4}\rightarrow\mathbb{R}P^{4}$ implies that there is a homotopy equivalence $\ast A(2p)\rightarrow A(2p)$. From (6.11) and the related cobordism, we have that the Kirby-Siebenmann invariant is $\operatorname{KS}(\ast A(2p))=\operatorname{KS}(\ast\mathbb{R}P^{4})=1\in\mathbb{Z}/2$ [16, Theorem 9.2] (see [16, Section 9.2] for a description of the construction of this invariant). Thus, the four-manifold $\ast A(2p)$ is non-smoothable for any odd $p\in\mathbb{Z}_{>0}$. $\square$